Friday, July 27, 2012

Insert tab A in ... which one is slot A?

Mark, Daniel, and Charles spent yesterday and today connecting (most of) the wires, fibers, hoses, cables, and tubes on ODI.  They have one or two more days before we turn the whole instrument on.  Everything is going well - other than the discovery that the chain that holds the cable (in an elongated loop in the bottom photo) is not the right size.  A replacement has been ordered, and this one will suffice for now.

Towards the end of Friday, a few of the subsystems were checked out.  The Dewar pumped down fine, and the shutter appears to operate correctly.  The ISP computer came up and is now monitoring the many temperature sensors within the instrument.  More next week.


No comments:

Post a Comment